# boxparampadding=1.5pt, padding-left=2pt, padding-right=2pt, height=7.5pt, background-color=lightgray, border-radius=2pt, border-right-width=1pt, border-bottom-width=1pt, border-color=gray ExploreSelf: Fostering User-driven Exploration and Reflection on Personal Challenges with Adaptive Guidance by Large Language Models

### v2

@@ Abstract.
Expressing stressful experiences in words is proven to improve mental and physical health, but individuals often disengage with writing interventions as they struggle to organize their thoughts and emotions. Reflective prompts have been used to provide direction, and large language models (LLMs) have demonstrated the potential to provide tailored guidance. Current systems often limit users’ flexibility to direct their reflections. We thus present ExploreSelf, an LLM-driven application designed to empower users to control their reflective journey. ExploreSelf allows users to receive adaptive support through dynamically generated questions. Through an exploratory study with 19 participants, we examine how participants explore and reflect on personal challenges using ExploreSelf. Our findings demonstrate that participants valued the flexible navigation of adaptive guidance to control their reflective journey, leading to deeper engagement and insight. Building on our findings, we discuss the implications of designing LLM-driven tools that facilitate user-driven and effective reflection of personal challenges.
@@ 1. Introduction
The generative capabilities of recent large language models (LLMs) have accelerated the advancement of intelligent writing support tools (Lee et al., 2024, 2022), which commonly perform natural language generation for directly modifying the content (e.g., (Kim et al., 2024b; Lee et al., 2022)) or providing advice for improving the content (e.g., (Evmenova et al., 2024)). However, research on LLM-driven writing tools largely focuses on creative writing, and only a handful of works have recently begun to study intelligent support for writing about the self (e.g., DiaryMate (Kim et al., 2024b)). Further, a growing body of research has investigated the LLMs’ potential for mental health support in the form of conventional chat interfaces (e.g., (Kim et al., 2024a; Seo et al., 2024; Kumar et al., 2024; Song et al., 2024)). Although LLM-driven, open-ended conversations are intuitively analogous to therapists’ counseling sessions, an agent who holds the therapist’s persona tends to lead the conversation and sometimes can result in unintentional reinforcement of negative narratives (Song et al., 2024; Inflection AI, 2023; Kim et al., 2024a). Meanwhile, research suggests that when individuals are able to make choices and exert control over their experiences, they engage more meaningfully in the process and develop a stronger sense of ownership and responsibility for their growth (Deci and Ryan, 2008).
In this work, we investigate ways to empower individuals to take greater control of the reflective process of their personal challenges while simultaneously leveraging the personalized and adaptive guidance offered by LLMs. To this aim, we designed and developed ExploreSelf, an LLM-infused interactive system for exploring and reflecting on personal challenges. ExploreSelf allows users to freely expand related themes from the initial narrative and drill down probing questions to reflect on aspects of themes in depth. To facilitate exploration, the system provides on-demand guidance, such as prompting keywords and comments tailored to their writing. The design of ExploreSelf was informed by formative interviews with nine mental health professionals, where their challenges and approaches to managing the counseling sessions inspired our interface design and the generative pipelines.
@@
The design and implementation of ExploreSelf, a novel LLM-driven web application that supports guided writing about breakdown topics regarding a personal issue, encouraging sensemaking and refreshment of thoughts on it. ExploreSelf’s design was informed by formative interviews () with clinical psychologists and professional licensed counselors.
@@
Implications for designing LLM-driven tools that facilitate user-driven and effective reflection of personal challenges.
@@ 2.1. Writing for Well-being
Research from psychology has indicated that writing about stressful events demands a significant level of perceived control, as it helps individuals organize, connect, and contextualize their thoughts and emotions (Andersson and Conley, 2008; Frazier et al., 2011). This sense of control is not only essential for initiating or sustaining a writing practice but is also closely tied to physical and psychological well-being, making it a core component of the writing process itself (Andersson and Conley, 2008). To support individuals in maintaining a sense of control, structured guidance has been studied as an effective tool to reduce cognitive load and alleviate the overwhelm that often accompanies reflection. Prompts and reflection frameworks are amongst the popularly studied tools for offering that sense of control, offering a sense of direction or goal to writers. In this way, individuals can navigate their thoughts and emotions in a more organized way, providing manageable steps to follow rather than leaving them to explore themselves entirely on their own. This approach supports sustained engagement by making the process more approachable and less overwhelmingDespite the benefits, many individuals fi. Previous research has explored various methods for designing such structure, including gratitude-focused writing (Fekete and Deichert, 2022), writing in a narrative structure (Danoff-Burg et al., 2010), using sentence stems (e.g., “The thing I am most worried about is…”), journaling with photographs (Nave et al., 2016), and mind maps (Danoff-Burg et al., 2010; Thompson, 2004). While these approaches offer helpful starting points, they are often provided in a fixed format, usually requiring the individual to stay on a structured course throughout, therefore they cannot respond to the writer’s thoughts and feelings that evolve in writing. This limits their ability to provide adaptive feedback or tailored suggestions to the individual’s unique emotional states and goals in writing. Recognizing this gap, recent works have explored more dynamic, interactive methods that provide personalized prompts, real-time feedback, and adaptive support (Miller, 2014; Wang et al., 2018; Jeong and Breazeal, 2016).
@@ 2.2. Technology-Mediated Writing for Well-being
Research suggests that when individuals are given the freedom to make choices and exercise control over their reflective processes, they tend to experience deeper engagement driven by intrinsic motivation (Patall et al., 2008). This autonomy fosters a stronger sense of ownership over their personal growth, increasing their responsibility for and investment in the outcomes of their reflection (Ryan and Deci, 2000; Deci and Ryan, 2008; Patall et al., 2008; Meng and Ma, 2015; Evans and Boucher, 2015). Recently in HCI, self-guided tools have been studied in the context of therapeutic intervention (e.g., (Sharma et al., 2023, 2024)), allowing individuals to engage with LLM-driven systems to reframe negative patterns of thinking. These systems aim to improve the accessibility and effectiveness of interventions by guiding users through well-established tasks, such as identifying cognitive distortions and reframing, with the LLM providing adaptive support throughout the cognitively and emotionally demanding process. However, while such predefined therapeutic modules provide focused support, in this case on correcting cognitive distortions, these systems may still keep the users from exploring the freely flowing thoughts and feelings of users and coming to a resolution or new self-insight on their own. In this work, we are interested in designing an LLM-driven writing for well-being activity that can preserve user agency throughout the reflection process while still providing adaptive support. Striking this balance between guidance and autonomy is crucial, and we turn to formative studies where we explore ways to integrate adaptive LLM support within the context of users’ own reflective journey.
@@ 3.1. Procedure and Analysis
Alias | Job title | Years of experience | Country of experience |
E1 | Licensed counselor | 16 years | South Korea |
E2 | Clinical psychologist | 26 years | South Korea |
E3 | Licensed counselor | 15 years | South Korea |
E4 | Licensed counselor | 17 years | South Korea |
E5 | Psychiatrist | 18 years | South Korea |
E6 | Clinical psychologist | 20 years | United States, United Kingdom |
E7 | Clinical psychologist | 11 years | United States |
E8 | Clinical psychologist | 5 years | United States |
E9 | Clinical psychologist | 8 years | United States |
We recruited nine experts (E1–9; see Table 1) through snowball sampling and internal network, who have extensive experience in clinical psychology and professional counseling. Participants comprised five clinical psychologists, three licensed therapists, and one psychiatrist. Five were based in South Korea, three in the United States, and one had experience working across the US and the UK. Participants had an average of 15.11 years (ranged 5 to 26) of experience in mental health care. Each interview lasted about one hour and was conducted in person or remotely, depending on the participants’ availability. We compensated participants with 100,000 KRW or 100 USD, depending on their residence.
The interviews covered participants’ strategies for helping clients navigate personal challenges, typical difficulties clients face, and potential roles they envision for LLM-driven systems in the therapeutic process. To elicit their feedback, we presented the experts with an example of a personal challenge and asked them to describe how they would approach such a case, including how they might guide the client through the process. Additionally, we presented a paper prototype of an LLM-driven guided writing where the system suggests related topics based on the initial problem narrative the user wrote. We asked for participants’ feedback on its potential utility and areas for improvement.
@@ 4.1. Design Rationales
DR3. Provide a tangible summary of the progress. In our formative study, professionals emphasized the importance of restating or summarizing client’s thoughts back to them, as this helps foster reflection by offering an objective perspective on the client’s experiences. They noted that while the summary might be subjective from the therapist’s viewpoint, it allows individuals to compare their internal thoughts with an external interpretation, serving as a valuable tool for stepping back and critically evaluating their journey. In line with this practice, we introduced the “AI Summary” feature, designed to offer users an objective summary of their reflections. The overview provides a synthesized narrative of the user’s themes, questions, and answers, reflecting the key points of their exploration.
@@ 4.2. System Design and User Interface
ExploreSelf is designed to assist users in navigating, exploring, and making sense of personal challenges through writing with structured and adaptive guidance by an LLM. The application consists of three primary phases of user engagement: (1) writing the initial narrative, (2) exploring the narrative, and (3) wrapping up and summarization. Each phase was carefully designed to encourage different aspects of self-reflection on and understanding of the personal challenges, helping users process their thoughts and emotions effectively. In the following, we describe each phase and interactions through a usage scenario: Jane is a retiree who looked forward to enjoying the freedom she earned after decades of hard work. She had plans to explore new hobbies and focus on self-development. However, her daughter asked Jane to care for her newborn during the daytime. Although Jane loves spending time with her grandson, she sometimes feels that her own life has become overshadowed by her new responsibilities. The freedom she anticipated seems out of reach, and she occasionally struggles with a sense of purposelessness, yearning for time to pursue her personal growth. Now that she turns to ExploreSelf to process this challenge.
@@ 4.4. Implementation
ExploreSelf consists of (1) a backend server that manages user data and LLM pipelines and (2) a frontend web interface, both written in TypeScript (Microsoft, 2024). The server runs on the Express.js (Foundation, 2024) framework, providing core functionalities via REST APIs and storing the user data in MongoDB (MongoDB, Inc., 2024). We implemented the LLM pipelines using LangChain.js (LangChain, 2024) incorporating OpenAI (OpenAI, 2024)’s GPT-4 ChatCompletion API. As LLM’s capabilities of dealing with non-English texts are inferior to equivalent English texts due to the inefficient tokenization (Petrov et al., 2023), we chose gpt-4o111https://openai.com/index/hello-gpt-4o/ as an underlying model, which involved an improved Korean tokenizer with state-of-the-art performance. The web interface was implemented using React.js (Meta, 2024).
@@ 5. User Study
We conducted an exploratory study to examine how individuals interact with ExploreSelf and how it may foster exploration and reflection on personal challenges. The study was conducted online via Zoom to ensure that participants could engage in the study in an environment of their choice, where they could feel most comfortable and secure. Conducting the study online allowed for greater flexibility in participation, enabling individuals to interact with ExploreSelf at their own pace and in their preferred settings, which is known to lead to more authentic and reflective responses (Schmader and Sedikides, 2018; Hoogesteyn et al., 2020). Our study protocol was approved by the public institutional review board of the Ministry of Health and Welfare of South Korea prior to recruitment.
@@ 5.1. Participants
Alias | Age | Gender | Occupation | Frequency of AI assistant use |
@@
P11 | 36 | Male | Full-time (Finance) | Tried once or twice |
@@
P14 | 65 | Female | Part-time (Elderly welfare) | At least once a day |
@@
P16 | 32 | Female | Full-time (Healthcare) | Tried once or twice |
P17 | 39 | Male | Private academy instructor | 4-6 times a week |
@@
P19 | 49 | Male | Full-time (Distribution) | Never |
We recruited 19 participants (P1–19; 10 female) by advertising the study on a local social community platform and online university communities in South Korea. Our inclusion criteria were adults who (1) are 19 years old or older; (2) are native Korean speakers; (3) are currently staying in Korea; (4) can join an online study in a private and comfortable place; and (5) are motivated to better understand their personal challenges. In particular, we ensured participants met criteria (2) and (3), (2) for them to be able to express their personal challenges in the most comfortable language as ExploreSelf was implemented in Korean. (3) was to make sure the participants were in an accessible proximity, allowing them to engage in the study from a private space of their choosing, where they could write without researchers physically present while still being accessible for the study. Please refer to Section 5.2 for more detail. The advertisement included a link to a screening survey that asked basic demographic information (i.e., age, gender, educational level, employment details, as well as background questions about prior experience with AI and personal endeavor to process personal challenges, if any.

### v3

@@ Abstract.
Expressing stressful experiences in words is proven to improve mental and physical health, but individuals often disengage with writing interventions as they struggle to organize their thoughts and emotions. Reflective prompts have been used to provide direction, and large language models (LLMs) have demonstrated the potential to provide tailored guidance. However, current systems often limit users’ flexibility to direct their reflections. We thus present ExploreSelf, an LLM-driven application designed to empower users to control their reflective journey, providing adaptive support through dynamically generated questions. Through an exploratory study with 19 participants, we examine how participants explore and reflect on personal challenges using ExploreSelf. Our findings demonstrate that participants valued the flexible navigation of adaptive guidance to control their reflective journey, leading to deeper engagement and insight. Building on our findings, we discuss the implications of designing LLM-driven tools that facilitate user-driven and effective reflection of personal challenges.
@@ 1. Introduction
The generative capabilities of recent large language models (LLMs) have accelerated the advancement of intelligent writing support tools (Lee et al., 2024a, 2022), which commonly perform natural language generation for directly modifying the content (e.g., (Kim et al., 2024b; Lee et al., 2022)) or providing advice for improving the content (e.g., (Evmenova et al., 2024)). However, research on LLM-driven writing tools largely focuses on creative writing, and only a handful of works have recently begun to study intelligent support for writing about the self (e.g., DiaryMate (Kim et al., 2024b)). Further, a growing body of research has investigated the LLMs’ potential for mental health support in the form of conventional chat interfaces (e.g., (Kim et al., 2024a; Seo et al., 2024; Kumar et al., 2024; Song et al., 2024)). Although LLM-driven, open-ended conversations are intuitively analogous to therapists’ counseling sessions, an agent who holds the therapist’s persona tends to lead the conversation and sometimes can result in unintentional reinforcement of negative narratives (Song et al., 2024; Inflection AI, 2023; Kim et al., 2024a). Meanwhile, research suggests that when individuals are able to make choices and exert control over their experiences, they engage more meaningfully in the process and develop a stronger sense of ownership and responsibility for their growth (Deci and Ryan, 2008).
In this work, we investigate ways to empower individuals to take greater control of the reflective process of their personal challenges while simultaneously leveraging the personalized and adaptive guidance offered by LLMs. Unlike traditional mental health services that focus on professional diagnosis and intervention, our goal was to design a system that enables individuals to explore their challenges at their own pace, without clinical oversight. To this aim, we designed and developed ExploreSelf (Figure 1), an LLM-infused interactive system for exploring and reflecting on personal challenges. ExploreSelf allows users to freely expand related themes from the initial narrative and drill down into probing questions to reflect on aspects of themes in depth. To facilitate exploration, the system provides on-demand guidance, such as prompting keywords and comments tailored to their writing. The design of ExploreSelf was informed by formative interviews with nine mental health professionals, where their challenges and approaches to managing the counseling sessions inspired our interface design and the generative pipelines.
@@
The design and implementation of ExploreSelf, a novel LLM-driven web application that supports guided writing about breakdown topics regarding a personal issue, encouraging sensemaking and refreshment of thoughts on it. ExploreSelf’s design was informed by formative interviews () with clinical psychologists and professional licensed counselors. The source code of ExploreSelf is publicly available at https://naver-ai.github.io/exploreself.
@@
Implications for designing LLM-guided tools that facilitate user-driven and effective reflection of personal challenges, such as balancing user autonomy with guided support, fostering long-term practice, and implementing the nuanced role and stance of AI to accommodate a diverse range of individuals.
@@ 2.1. Writing for Well-being
Research from psychology has indicated that writing about stressful events demands a significant level of perceived control, as it helps individuals organize, connect, and contextualize their thoughts and emotions (Andersson and Conley, 2008; Frazier et al., 2011). This sense of control is not only essential for initiating or sustaining a writing practice but is also closely tied to physical and psychological well-being, making it a core component of the writing process itself (Andersson and Conley, 2008). To support individuals in maintaining a sense of control, structured guidance has been studied as an effective tool to reduce cognitive load and alleviate the overwhelm that often accompanies reflection. Prompts and reflection frameworks are amongst the popularly studied tools for offering that sense of control, offering a sense of direction or goal to writers. In this way, individuals can navigate their thoughts and emotions in a more organized way, providing manageable steps to follow rather than leaving them to explore themselves entirely on their own. This approach supports sustained engagement by making the process more approachable and less overwhelming (Cleary and Zimmerman, 2004; Bandura, 1991). Previous research has explored various methods for designing such structure, including gratitude-focused writing (Fekete and Deichert, 2022), writing in a narrative structure (Danoff-Burg et al., 2010), using sentence stems (e.g., “The thing I am most worried about is…”), journaling with photographs (Nave et al., 2016), and mind maps (Danoff-Burg et al., 2010; Thompson, 2004). While these approaches offer helpful starting points, they are often provided in a fixed format, usually requiring the individual to stay on a structured course throughout, therefore they cannot respond to the writer’s thoughts and feelings that evolve in writing. This limits their ability to provide adaptive feedback or tailored suggestions to the individual’s unique emotional states and goals in writing. Recognizing this gap, recent works have explored more interactive methods that provide personalized prompts, real-time feedback, and adaptive support (Miller, 2014; Wang et al., 2018; Jeong and Breazeal, 2016).
@@ 2.2. Technology-Mediated Writing for Well-being
Research suggests that when individuals are given the freedom to make choices and exercise control over their reflective processes, they tend to experience deeper engagement driven by intrinsic motivation (Patall et al., 2008). This autonomy fosters a stronger sense of ownership over their personal growth, increasing their responsibility for and investment in the outcomes of their reflection (Ryan and Deci, 2000; Deci and Ryan, 2008; Patall et al., 2008; Meng and Ma, 2015; Evans and Boucher, 2015). Recently in HCI, self-guided tools have been studied in the context of therapeutic intervention (e.g., (Sharma et al., 2023, 2024)), allowing individuals to engage with LLM-driven systems to reframe negative patterns of thinking. These systems aim to improve the accessibility and effectiveness of interventions by guiding users through well-established tasks, such as identifying cognitive distortions and reframing, with the LLM providing adaptive support throughout the cognitively and emotionally demanding process. However, while such predefined therapeutic modules provide focused support, in this case on correcting cognitive distortions, these systems may still keep the users from exploring the freely flowing thoughts and feelings of users and coming to a resolution or new self-insight on their own. In this work, we are interested in designing an LLM-driven writing for well-being activity that can preserve user agency throughout the reflection process while still providing adaptive support. Striking this balance between guidance and autonomy is crucial, and we turn to formative studies where we explore ways to integrate adaptive LLM support within the context of users’ reflective journeys.
@@ 3.1. Procedure and Analysis
| Alias | Gender | Job title |
Years of
experience |
Country of
experience |
| E1 | Female | Licensed counselor | 16 years | South Korea |
| E2 | Female | Clinical psychologist | 26 years | South Korea |
| E3 | Female | Licensed counselor | 15 years | South Korea |
| E4 | Female | Licensed counselor | 17 years | South Korea |
| E5 | Female | Psychiatrist | 18 years | South Korea |
| E6 | Female | Clinical psychologist | 20 years |
United
States,
United Kingdom |
| E7 | Female | Clinical psychologist | 11 years | United States |
| E8 | Male | Clinical psychologist | 5 years | United States |
| E9 | Female | Clinical psychologist | 8 years | United States |
We recruited nine experts (E1–9, eight females and one male; see Table 1) through snowball sampling and internal network, who have extensive experience in clinical psychology and professional counseling. We initially planned to recruit experts in South Korea as our system was intended to be evaluated with the Korean population. However, the prolonged medical crisis in South Korea in 2024 made it difficult to recruit clinical psychologists locally. To supplement expert feedback from clinical psychologists, we recruited four clinical psychologists from outside South Korea through our internal network. The final nine participants comprised five clinical psychologists, three licensed therapists, and one psychiatrist. Five were based in South Korea, three in the United States, and one had experience working across the United States and the United Kingdom. Participants had an average of 15.11 years (ranged 5 to 26) of experience in mental health care. Each interview lasted about one hour and was conducted in person or remotely, depending on the participants’ availability. We compensated participants with 100,000 KRW or 100 USD, depending on their residence.
The interviews covered participants’ strategies for helping clients navigate personal challenges, typical difficulties clients face, and potential roles they envision for LLM-driven systems in the therapeutic process. Since the experts have cared for various client populations in different settings, we asked them to share insights applicable to the general public. To elicit their feedback, we also presented the experts with an example personal challenge narrative (see Appendix A.1) and asked them to describe how they would approach such a case, including how they might guide the client through the process. Three researchers devised the example narrative during a brainstorming session, focusing on creating culturally neutral challenges. Additionally, we presented a paper prototype of an LLM-driven guided writing where the system suggests related topics based on the same example personal challenge narrative. We asked for participants’ feedback on its potential utility and areas for improvement.
@@ 4.1. Design Rationales
DR3. Provide a tangible summary of the progress. In our formative study, professionals emphasized the importance of restating or summarizing client’s thoughts back to them, as this helps foster reflection by offering an objective perspective on the client’s experiences. They noted that while the summary might be subjective from the therapist’s viewpoint, it allows individuals to compare their internal thoughts with an external interpretation, serving as a valuable tool for stepping back and critically evaluating their journey. In line with this practice, we introduced the “AI Summary” feature, designed to offer users an objective summary of their reflections, distilled from the user’s themes, questions, and answers, reflecting the key points of their exploration.
@@ 4.2. System Design and User Interface
ExploreSelf is designed to assist users in navigating, exploring, and making sense of personal challenges through writing with structured and adaptive guidance by an LLM. The application consists of three primary phases of user engagement: (1) Writing the initial narrative (Figure 2), (2) Exploring the narrative (Figure 3), and (3) Wrapping up and summarization (Figure 4). Each phase was carefully designed to encourage different aspects of self-reflection on and understanding of the personal challenges, helping users process their thoughts and emotions effectively. In the following, we describe each phase and interactions through a usage scenario: Jane is a retiree who looked forward to enjoying the freedom she earned after decades of hard work. She had plans to explore new hobbies and focus on self-development. However, her daughter asked Jane to care for her newborn during the daytime. Although Jane loves spending time with her grandson, she sometimes feels that her own life has become overshadowed by her new responsibilities. The freedom she anticipated seems out of reach, and she occasionally struggles with a sense of purposelessness, yearning for time to pursue her personal growth. Now that she turns to ExploreSelf to process this challenge.
@@ 4.4. Implementation
ExploreSelf consists of (1) a backend server that manages user data and LLM pipelines and (2) a frontend web interface, both written in TypeScript (Microsoft, 2024). The server runs on the Express.js (Foundation, 2024) framework, providing core functionalities via REST APIs and storing the user data in MongoDB (MongoDB, Inc., 2024). We implemented the LLM pipelines using LangChain.js (LangChain, 2024) incorporating OpenAI (OpenAI, 2024b)’s GPT-4 ChatCompletion API. As LLM’s capabilities of dealing with non-English texts are inferior to equivalent English texts due to the inefficient tokenization (Petrov et al., 2023), we chose gpt-4o111https://openai.com/index/hello-gpt-4o/ as an underlying model, which involved an improved Korean tokenizer with state-of-the-art performance. The web interface was implemented using React (Meta, 2024).
@@ 5. User Study
We conducted an exploratory study to examine how individuals interact with ExploreSelf and how it may foster exploration and reflection on personal challenges. As part of the study, participants explored their own challenges using ExploreSelf on their computers. The study was conducted online via Zoom to allow participants to engage from a comfortable and secure environment of their choice. This online approach enabled participants to interact with ExploreSelf at their own pace and in their preferred settings, which has been shown to foster more authentic and reflective responses (Schmader and Sedikides, 2018; Hoogesteyn et al., 2020). We refined both the system and the study protocol through three pilot sessions with Korean adults. For example, we incorporated the Pinned Themes feature, enabling users to pin and revisit previously generated themes. We also revised the wording used in the interface and the tutorial procedure to clarify that the system is not intended to provide solutions or counseling but to support thematic exploration. The study protocol was approved by the public institutional review board of the Ministry of Health and Welfare of South Korea.
@@ 5.1. Participants
| Alias | Age | Gender | Occupation |
Frequency of
AI assistant use |
@@
| P11 | 36 | Male |
Full-time
Tried once or twice |
@@
| P14 | 65 | Female |
Part-time
(Elderly welfare) |
At least once a day |
@@
| P16 | 32 | Female |
Full-time
Tried once or twice |
| P17 | 39 | Male |
Private academy
instructor |
4-6 times a week |
@@
| P19 | 49 | Male |
Full-time
Never |
We recruited 19 participants (P1–19; 10 female) by advertising the study on a local social community platform and online university communities in South Korea. Our inclusion criteria were adults who (1) are 19 years old or older; (2) are native Korean speakers; (3) are currently staying in Korea; (4) can join an online study in a private and comfortable place; and (5) are motivated to better understand their personal challenges. In particular, we ensured participants met criteria (2) and (3), (2) for them to be able to express their personal challenges in the most comfortable language as ExploreSelf was implemented in Korean. (3) was to make sure the participants were in an accessible proximity, allowing them to engage in the study from a private space of their choosing, where they could write without researchers physically present while still being accessible for the study. Please refer to Section 5.2 for more detail. The advertisement included a link to a screening survey that asked basic demographic information (i.e., age, gender, educational level, employment details, as well as background questions about prior experience with AI and personal endeavor to process personal challenges, if any. We did not screen participants for the presence of mental illness, as ExploreSelf was designed to support the general population’s mental wellbeing rather than serve as a psychotherapeutic intervention.